.0 W/m2-K, including the effec An anaerobic digester cover has an overall heat transfer coefficient U of 1.0 W/m2.K, including the effects of convection and radiation. The cover has a surface area of 325 m2. The temperature inside the digester is 33°C and the outside air temperature is 0.0°C. What is the rate of heat loss from the digester through the cover? Report your answer in kW, to the nearest 0.1 kW, without entering the units.
Answer #1

Q= UA\DeltaT

U:OVER ALL HEAT TRANSFERE COEFFI IENT =1W/m2K

A; SURFACE AEA =325m2

\DeltaT= temperature difference=33-0=33k

Q=1*325*33=10.725KW